It is claimed that automobiles are driven on average more than 20,000 kilometers per year. To test this claim, 100 randomly selected automobile owners are asked to keep a record of the kilometers they travel. Would you agree with this claim if the random sample showed an average of 23,500 kilometers and a standard deviation of 3900 kilometers?
1
Expert's answer
2020-05-18T19:42:24-0400

sample size n=100

sample mean x̄=23500

standard deviation σ=3900

the null hypothesis is:

H0: μ = 20000

the alternative hypothesis is:

H1: μ > 20000

z=(x̄-μ)/(σ/"\sqrt n" )=(23500-20000)/(3900/10)=8.97

P(Z>8.97)=1-P(Z<8.97)=1-1=0

P-value is approximately 0, so we reject the null hypothesis and conclude that automobiles are driven on average more than 20,000 kilometers per year